OpenStudy (anonymous):

1+sin^x=3sinxcosx than tan x=1 or1/2

OpenStudy (raden):

sin^x ?

OpenStudy (shubhamsrg):

perhaps you meant sin^2 x ? please confirm

OpenStudy (anonymous):

1+sin^2x=3sinxcosx cos^2x+sin^2x+sin^2x=3sinxcosx (cos^2x+sin^2x=1) cos^2 x+2 sin^2 x=3sinxcosx Divide by cos^2 x, 1+2 tan^2 x=3 tan x 2 tan^2 x -3 tan x+1=0 2tan^2 x-2tanx-tanx+1=0 2tanx(tanx-1)-1(tanx-1)=0 (tanx-1)(2tanx-1)=0 tanx-1=0,tanx=1 or2tanx-1=0,2tanx=1,tanx=1/2